A school sold 200 tickets. Adult tickets cost 6 dollars each and student tickets cost 4 dollars each. Total sales were 940 dollars. how many of each ticket were sold?
Answer by checkley75(3666)

A+S=200 OR A=200-S
6A+4S=940 NOW SUBSTITUTE (200-S) FOR A & SOLVE FOR S
6(200-S)+4S=940
1200-6S+4S=940
-2S=940-1200
-2S=-260
S=-260/-2
S=130 TICKETS SOLD TO STUDENTS THEN
200-130=70 ADULT TICKETS SOLD.
PROOF
6*70+4*130=940
420+520=940
940=940